OpenStudy (anonymous):

In 25 in which _?_ must be raised in order to produce 25.

OpenStudy (anonymous):

5^2..

OpenStudy (anonymous):

e^x=25

OpenStudy (anonymous):

Ln is there..

OpenStudy (anonymous):



OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

\[5^2 = 25 \implies Log_5 25= 2\]

OpenStudy (anonymous):

so 5^2=25

OpenStudy (anonymous):